# The Picard group of the Baily-Borel compactification of the moduli space of polarized K3 surfaces

### v1

@@ 4.1 Projection to ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M})$
and $G_{r}^{h}(1,1)=\binom{r-3+h}{h}$. Then we can define the differential operator
@@
where the coordinate $Z=\left(\begin{array}[]{cc}z_{1}&z_{2}\\ z_{3}&z_{4}\end{array}\right)\in\mathbb{H}_{2}$. For the Theta series, we have the following facts.
@@
Suppose $h>0$. Let $\{H_{i}\}$ be an orthonormal basis of the space of harmonic forms of degree $h$. For $(z,z^{\prime})\in\mathbb{H}\times\mathbb{H}\subseteq\mathbb{H}_{2}$, we have
@@
$(\partial_{h}\Theta^{(2)}_{M,1})(z,z^{\prime})=C\sum\limits_{i}\Theta^{(1)}_{M,H_{i}}(z)\otimes\Theta^{(1)}_{M,\overline{H}_{i}}(z^{\prime})$ (4.2)
@@
For the first assertion, this follows from the definition of the Theta series. The second assertion then follows from [25, Proposition 2.7].
Let $k=\frac{r}{2}+h$ for some non-negative integer $h$. When $h=0$, assume further $r>4$. As in [25], setting $\vartheta_{M,k}:=\partial_{h}\Theta_{\mathbf{Gen}(M)}^{(2)}$, we can define a Theta lifting map
@@
by sending $f$ to
$\int_{{\rm SL}_{2}(\mathbb{Z})\backslash\mathbb{H}}\langle f(z),\vartheta_{M,k}(z,-\overline{z^{\prime}})\rangle y^{k}\frac{\mathrm{d}x\mathrm{d}y}{y^{2}}$
as a function in $z^{\prime}$. Then one has
@@
Let $k,M$ be as above. The linear map $\Psi$ is diagonalizable and surjective onto ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M})$. In particular, there is a decomposition
@@
The proof is essentially [25, Proposition 6.2]. According to Lemma 4.1, we have
$\Psi(f)(z^{\prime})$ $=\int_{{\rm SL}_{2}(\mathbb{Z})\backslash\mathbb{H}}\left(\sum\limits_{L\in\mathbf{Gen}(M)}\sum\limits_{\sigma\in\mathrm{Iso}(G_{M},G_{L})}C_{L,\sigma}\langle f(z),\sigma^{\ast}\partial_{h}\Theta_{L,1}^{(2)}(z,-\overline{z^{\prime}})\rangle\right)y^{k}\frac{\mathrm{d}x\mathrm{d}y}{y^{2}}$
$=C\sum\limits_{i}\sum\limits_{L\in\mathbf{Gen}(M)}\sum\limits_{\sigma\in\mathrm{Iso}(G_{M},G_{L})}C_{L,\sigma}\langle f,\sigma^{\ast}\Theta_{L,H_{i}}^{(1)}\rangle_{\rm Pet}\cdot\sigma^{\ast}\Theta_{L,H_{i}}^{(1)}(z^{\prime})$
where the constant $C$ and the harmonic polynomials $\{P_{i}\}$ are those in Lemma 4.1, and
@@
is given in (2.5). When $h>0$, as $\Theta_{L,H_{i}}^{(1)}$ is cuspidal, we have $\Psi(f)\in{\mathrm{Cusp}}^{\theta}_{k}(\rho_{M})$. When $h=0$, we have all $H_{i}=1$. Note that $\Theta_{\mathbf{Gen}(M)}^{(1)}=\mathbf{E}^{(1)}_{k,M}$ (by Theorem 2.5) is orthogonal to $f$ under the Petersson inner product, one can rewrite $\Psi(f)$ as
$\Psi(f)(z^{\prime})=\sum\limits_{L\in\mathbf{Gen}(M)}\sum\limits_{\sigma\in\mathrm{Iso}(G_{M},G_{L})}C_{L,\sigma}\langle f,\sigma^{*}\Theta_{L,1}-\Theta^{(1)}_{\mathbf{Gen}(M)}\rangle_{\rm Pet}\cdot(\sigma^{*}\Theta_{L,1}-\Theta^{(1)}_{\mathbf{Gen}(M)}).$
@@
Thus a statement like "${\mathrm{Cusp}}={\mathrm{Cusp}}^{\theta}$" is equivalent to "$\ker\Phi=0$".
4.2 $\Psi$ as Hecke operators
Let $\alpha\in\mathbb{Z}^{>0}$. Let $K_{\alpha}(z,z^{\prime}):\mathbb{H}\times\mathbb{H}\to\mathbb{C}[G_{M}]\otimes\mathbb{C}[G_{M}]$ be given by
$\sum\limits_{\gamma\in G_{M}}\sum\limits_{\tilde{g}=(g,\phi_{g}(z))\in\widetilde{\mathbf{Y}}_{\alpha^{2}}}\frac{\phi_{g}(z)^{-2k}}{(z^{\prime}+g\cdot z)^{k}}\rho_{M}(\widetilde{g})^{-1}(\mathfrak{e}_{\gamma})\otimes\mathfrak{e}_{\gamma}.$ (4.4)
@@ 4.2 $\Psi$ as Hecke operators
As a function of $z$, one can check $\mathbf{T}_{\alpha^{2}}(K_{\alpha}(z,z^{\prime}))=\alpha^{2k-2}K_{\alpha}(z,z^{\prime})$. As shown in [25, Proposition 3.7] (see also Appendix C for the odd signature case), it can be viewed as the kernel function of $\mathbf{T}_{\alpha^{2}}$ up to a scalar, i.e. it satisfies
@@
Assume that $r={\rm rank}(M)>6$, and let $k=\frac{r}{2}+h$ for some non-negative integer $h$. For any $f\in{\mathrm{Cusp}}_{k}(\rho_{M})$, we have
@@
Let $\widetilde{C}:=h!\binom{r-3+h}{h}\binom{k-1}{h}$. For $\tau=(z,z^{\prime})\in\mathbb{H}\times\mathbb{H}\subseteq\mathbb{H}_{2}$, we claim that
$\vartheta_{M,k}(\tau)=\partial_{h}\Theta^{(2)}_{\mathbf{Gen}(M)}=\begin{cases}\displaystyle\mathbf{E}_{k,M}^{(1)}(z)\otimes\mathbf{E}_{k,M}^{(1)}(z^{\prime})+\frac{\mathbf{e}(\mathrm{sign}(M)/8)}{|G_{M}|^{\frac{1}{2}}}\sum\limits_{\alpha=1}^{\infty}K_{\alpha}(z,z^{\prime})&\hbox{if $h=0$}\\ \\ \displaystyle\widetilde{C}\frac{\mathbf{e}(\mathrm{sign}(M)/8)}{|G_{M}|^{\frac{1}{2}}}\sum\limits_{\alpha=1}^{\infty}\alpha^{h}K_{\alpha}(z,z^{\prime})&\hbox{if $h>0$}\end{cases}$ (4.7)
@@
$\int_{{\rm SL}_{2}(\mathbb{Z})\backslash\mathbb{H}}\langle f(z),\mathbf{E}_{k,M}^{(1)}(z)\otimes\mathbf{E}_{k,M}^{(1)}(-\overline{z^{\prime}}))\rangle y^{k}\frac{\mathrm{d}x\mathrm{d}y}{y^{2}}=\langle f(z),\mathbf{E}_{k,M}^{(1)}(z)\rangle_{\rm Pet}\cdot\mathbf{E}_{k,M}^{(1)}(z^{\prime})=0$ (4.8)
@@
$\Psi(f)$ $=\int_{{\rm SL}_{2}(\mathbb{Z})\backslash\mathbb{H}}\langle f(z),\vartheta_{M,k}(z,-\overline{z^{\prime}})\rangle y^{k}\frac{\mathrm{d}x\mathrm{d}y}{y^{2}}$
@@
$=\widetilde{C}\frac{\mathbf{e}(-\mathrm{sign}(M)/8)}{|G_{M}|^{\frac{1}{2}}}C^{-1}\sum\limits_{\alpha=1}^{\infty}\frac{1}{\alpha^{2k-2-h}}\mathbf{T}_{\alpha^{2}}(f),$
@@
To prove the claim, let us first collect some facts.
  1. $(i)$
    By Theorem 2.5, we have
@@ 2. $(ii)$
  3. $(iii)$
    As shown in the proof of [25, Theorem 4.7], recall $k=\frac{r}{2}+h$, we have
@@ 3. $(iii)$
Now we are ready to derive the formula. For simplicity, let us first assume $h>0$. Then by (4.11), (4.10) and (B.2), we have
@@
The argument is essentially the same (see [25, Theorem 4.6]) and we omit the details.
4.3 Injectivity of the projection
@@
Assume that $M$ is a negative definite even lattice of level $N$ and rank $r>6$. Furthermore, assume that for all primes $p$, $M\otimes\mathbb{Z}_{p}$ splits a hyperbolic plane. For $k\geq r/2$, we have
@@
As mentioned before, it suffices to show that the linear map $\Psi$ is injective. When $\mathrm{sign}(M)$ is even, this is exactly [25, Theorem 6.7]. When $\mathrm{sign}(M)$ is odd, the argument is very similar and we may sketch the proof for the ease of readers.
  1. $(i)$
@@
    $\Psi(f)$ $=C^{\prime}\sum\limits_{\alpha=1}^{\infty}\frac{\mathbf{T}_{\alpha^{2}}}{\alpha^{2k-2-h}}(f)=C^{\prime}\left(\sum_{\begin{subarray}{c}\alpha\geq 1\\ \gcd(\alpha,N)=1\end{subarray}}\frac{\mathbf{T}_{\alpha^{2}}}{\alpha^{2k-2-h}}\right)\left(\prod\limits_{p\mid N}\sum_{n=0}^{\infty}\frac{\mathbf{T}_{p^{2n}}}{p^{n(2k-2-h)}}\right)(f)$
@@
    Therefore the operator $\sum\limits_{\begin{subarray}{c}\alpha\geq 1\\ \gcd(\alpha,N)=1\end{subarray}}\frac{\mathbf{T}_{\alpha^{2}}}{\alpha^{2k-2-h}}$ is invertible, which implies $\prod\limits_{p\mid N}\sum\limits_{n=1}^{\infty}\frac{\mathbf{T}_{p^{2n}}}{p^{n(2k-2-h)}}(f)=0$.
  2. $(ii)$
    Let $n:=\prod\limits_{p\mid N}p$ be the radical of $N$. By [25, Lemma 6.6], according to our assumption on $M$, there exists a sublattice $L\subseteq M$ such that $G_{L}\cong G_{M}\oplus G^{\prime}$, where $G^{\prime}=\left(\begin{array}[]{cc}0&\frac{1}{n}\\ \frac{1}{n}&0\end{array}\right)\mod\mathbb{Z}$ under a basis $\{x_{1},x_{2}\}$ and $nx_{i}=0$. Then the cyclic subgroup $H=\langle x_{1}\rangle$ is an isotropic subgroup of $G_{L}$ and $H^{\perp}/H\cong G_{M}$.
  3. $(iii)$
@@
  4. $(iv)$
    According to [25, Lemma 6.3], also $\prod\limits_{p\mid N}\sum\limits_{n=1}^{\infty}\frac{\mathbf{T}_{p^{2n}}}{p^{n(2k-2-h)}}\uparrow_{H}(f)=0$. Let $v=v_{x_{1}+\gamma,x_{2}+\gamma,S}$ and then
@@ 4. $(iv)$
    where we used Proposition 3.7 for the last equality. Since $\gamma$ is arbitrary, this implies that $f=0$.
@@
There is a well-established criterion for determining whether $M\otimes\mathbb{Z}_{p}$ splits a hyperbolic plane, as mentioned in [27]. For example, this occurs if ${\rm rank}_{p}(G_{M})<r-2$. In practical terms, this condition is not overly restrictive when $r$ is sufficiently large, as ${\rm rank}_{p}(G_{M})\leq r$ always holds.
5 Heegner divisors on Shimura varieties of orthogonal type
Throughout this section, we let $M$ be an even lattice of signature $(2,n)$ under a pairing $\langle$-,$\rangle$ - , - ⟩ and $G_{M}=M^{\vee}/M$. For a normal variety $X$, ${\rm Cl}_{\mathbb{Q}}(X)$ will denote the class group of Weil divisors on $X$ with rational coefficients.
5.1 Shimura variety attached to a lattice

### v2

@@ 4.1. Projection to ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M})$
and $G_{r}^{h}(1,1)=\binom{r-3+h}{h}$. The differential operator is then defined by:
@@
where the coordinate $Z=\left(\begin{array}[]{cc}z_{1}&z_{2}\\ z_{2}&z_{4}\end{array}\right)\in\mathbb{H}_{2}$. For the theta series, we have the following facts.
@@
Suppose $h>0$. Let $\{H_{i}\}$ be an orthonormal basis for the space of harmonic polynomials of degree $h$. For $(z,z^{\prime})\in\mathbb{H}\times\mathbb{H}\subseteq\mathbb{H}_{2}$, we have
@@
$(\partial_{h}\Theta^{(2)}_{M,1})(z,z^{\prime})=C\sum\limits_{i}\Theta^{(1)}_{M,F_{i}}(z)\otimes\Theta^{(1)}_{M,\overline{F}_{i}}(z^{\prime})$ (4.2)
@@
For the first assertion, this follows from the definition of the theta series. The second assertion then follows from [Ma24, Proposition 2.7].
Let $k=\frac{r}{2}+h$ for some non-negative integer $h$. When $h=0$, assume further $r>4$. As in [Ma24], setting $\vartheta_{M,k}\coloneqq\partial_{h}\Theta_{\mathbf{Gen}(M)}^{(2)}$, we can define a theta lifting map
@@
by
$\Psi(f)(z^{\prime})=\int_{{\rm SL}_{2}(\mathbb{Z})\backslash\mathbb{H}}\langle f(z),\vartheta_{M,k}(z,-\overline{z^{\prime}})\rangle y^{k}\frac{\mathrm{d}x\mathrm{d}y}{y^{2}}$
regarded as a function in $z^{\prime}$. Then we have
@@
Let $k,M$ be as above. The linear map $\Psi$ is diagonalizable and surjective onto ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M})$. Consequently, there is a decomposition
@@
The proof is essentially [Ma24, Proposition 6.2]. According to Lemma 4.1, $\Psi(f)(z^{\prime})$ is equal to
$\int_{{\rm SL}_{2}(\mathbb{Z})\backslash\mathbb{H}}\left(\sum\limits_{L\in\mathbf{Gen}(M)}\sum\limits_{\sigma\in\mathrm{Iso}(G_{M},G_{L})}C_{L,\sigma}\langle f(z),\sigma^{\ast}\partial_{h}\Theta_{L,1}^{(2)}(z,-\overline{z^{\prime}})\rangle\right)y^{k}\frac{\mathrm{d}x\mathrm{d}y}{y^{2}}$
$=$ $\sum\limits_{i}\sum\limits_{L\in\mathbf{Gen}(M)}\sum\limits_{\sigma\in\mathrm{Iso}(G_{M},G_{L})}C_{L,\sigma}\langle f,\sigma^{\ast}\Theta_{L,F_{i}}^{(1)}\rangle_{\rm Pet}\cdot\sigma^{\ast}\Theta_{L,F_{i}}^{(1)}(z^{\prime})$
where the constant $C$ and the harmonic polynomials $\{H_{i}\}$ are those in Lemma 4.1, and
@@
is given in (2.5). When $h>0$, as $\Theta_{L,F_{i}}^{(1)}$ is cuspidal, we have $\Psi(f)\in{\mathrm{Cusp}}^{\theta}_{k}(\rho_{M})$. When $h=0$, we have all $H_{i}=1$. Note that $\Theta_{\mathbf{Gen}(M)}^{(1)}=\mathbf{E}^{(1)}_{k,M}$ (by Theorem 2.5) is orthogonal to $f$ under the Petersson inner product, one can rewrite $\Psi(f)$ as
$\sum\limits_{L\in\mathbf{Gen}(M)}\sum\limits_{\sigma\in\mathrm{Iso}(G_{M},G_{L})}C_{L,\sigma}\langle f,\sigma^{*}\Theta_{L,1}-\Theta^{(1)}_{\mathbf{Gen}(M)}\rangle_{\rm Pet}\cdot(\sigma^{*}\Theta_{L,1}-\Theta^{(1)}_{\mathbf{Gen}(M)}).$
@@
$f(v)=\sum_{i\in I}\langle v,v_{i}\rangle\cdot v_{i}$
is diagonalizable and surjective onto $\mathrm{Span}\{v_{i}:i\in I\}$.
Thus, a statement like "${\mathrm{Cusp}}={\mathrm{Cusp}}^{\theta}$" is equivalent to "$\ker\Psi=0$".
4.2. $\Psi$ as Hecke operators
Let $\alpha\in\mathbb{Z}_{>0}$. Define $K_{\alpha}(z,z^{\prime}):\mathbb{H}\times\mathbb{H}\to\mathbb{C}[G_{M}]\otimes\mathbb{C}[G_{M}]$ by
$K_{\alpha}(z,z^{\prime})=\sum\limits_{\gamma\in G_{M}}\sum\limits_{\tilde{g}=(g,\phi_{g}(z))\in\widetilde{\mathbf{Y}}_{\alpha^{2}}}\frac{\phi_{g}(z)^{-2k}}{(z^{\prime}+g\cdot z)^{k}}\rho_{M}(\widetilde{g})^{-1}(\mathfrak{e}_{\gamma})\otimes\mathfrak{e}_{\gamma}.$ (4.4)
@@ 4.2. $\Psi$ as Hecke operators
As a function of $z$, one can check $\mathbf{T}_{\alpha^{2}}(K_{\alpha}(z,z^{\prime}))=\alpha^{2k-2}K_{\alpha}(z,z^{\prime})$. As shown in [Ma24, Proposition 3.7] (see also Appendix C for the odd signature case), it can be viewed as the kernel function of $\mathbf{T}_{\alpha^{2}}$ up to a scalar, i.e. it satisfies
@@
Assume $r={\rm rank}(M)>6$, and let $k=\frac{r}{2}+h$ for some non-negative integer $h$. For any $f\in{\mathrm{Cusp}}_{k}(\rho_{M})$, we have
@@
Let $\widetilde{C}\coloneqq h!\binom{r-3+h}{h}\binom{k-1}{h}$. For $\tau=(z,z^{\prime})\in\mathbb{H}\times\mathbb{H}\subseteq\mathbb{H}_{2}$, we claim that
$\vartheta_{M,k}(\tau)=\partial_{h}\Theta^{(2)}_{\mathbf{Gen}(M)}\\ =\begin{cases}\displaystyle\mathbf{E}_{k,M}^{(1)}(z)\otimes\mathbf{E}_{k,M}^{(1)}(z^{\prime})+\frac{\mathbf{e}(\mathrm{sign}(M)/8)}{|G_{M}|^{\frac{1}{2}}}\sum\limits_{\alpha=1}^{\infty}K_{\alpha}(z,z^{\prime})&\hbox{if $h=0$}\\ \\ \displaystyle\widetilde{C}\frac{\mathbf{e}(\mathrm{sign}(M)/8)}{|G_{M}|^{\frac{1}{2}}}\sum\limits_{\alpha=1}^{\infty}\alpha^{h}K_{\alpha}(z,z^{\prime})&\hbox{if $h>0$}\end{cases}$ (4.7)
@@
$\int_{{\rm SL}_{2}(\mathbb{Z})\backslash\mathbb{H}}\langle f(z),\mathbf{E}_{k,M}^{(1)}(z)\otimes\mathbf{E}_{k,M}^{(1)}(-\overline{z^{\prime}}))\rangle y^{k}\frac{\mathrm{d}x\mathrm{d}y}{y^{2}}\\ =\langle f(z),\mathbf{E}_{k,M}^{(1)}(z)\rangle_{\rm Pet}\cdot\mathbf{E}_{k,M}^{(1)}(z^{\prime})=0$
@@
$\Psi(f)(z^{\prime})$ $=\int_{{\rm SL}_{2}(\mathbb{Z})\backslash\mathbb{H}}\langle f(z),\vartheta_{M,k}(z,-\overline{z^{\prime}})\rangle y^{k}\frac{\mathrm{d}x\mathrm{d}y}{y^{2}}$
@@
$=\widetilde{C}\frac{\mathbf{e}(-\mathrm{sign}(M)/8)}{|G_{M}|^{\frac{1}{2}}}C^{-1}\sum\limits_{\alpha=1}^{\infty}\frac{1}{\alpha^{2k-2-h}}\mathbf{T}_{\alpha^{2}}(f)(z^{\prime}),$
@@
To prove the claim, let us first collect some facts. By Theorem 2.5, we have
@@
As shown in the proof of [Ma24, Thm. 4.7], recall $k=\frac{r}{2}+h$, we have
@@
Now we are ready to derive the formula. Let us first assume $h>0$. By (4.10), (4.9) and (4.8), we then have
@@
The argument is essentially the same (see [Ma24, Thm. 4.6]) and we omit the details.
4.3. Injectivity of the projection
@@
Assume that $M$ is a positive definite even lattice of level $N$ and rank $r>6$. Furthermore, assume that for all primes $p$, $M\otimes\mathbb{Z}_{p}$ splits a hyperbolic plane. For $k\geq r/2$, we have
@@
As mentioned before, it suffices to show that the linear map $\Psi$ is injective. When $\mathrm{sign}(M)$ is even, this is exactly [Ma24, Thm. 6.7]. When $\mathrm{sign}(M)$ is odd, the argument is very similar, and we may sketch the proof for the ease of readers.
@@
$\Psi(f)$ $=C^{\prime}\sum\limits_{\alpha=1}^{\infty}\frac{\mathbf{T}_{\alpha^{2}}}{\alpha^{2k-2-h}}(f)$
$=C^{\prime}\left(\sum_{\begin{subarray}{c}\alpha\geq 1\\ \gcd(\alpha,N)=1\end{subarray}}\frac{\mathbf{T}_{\alpha^{2}}}{\alpha^{2k-2-h}}\right)\left(\prod\limits_{p\mid N}\sum_{n=0}^{\infty}\frac{\mathbf{T}_{p^{2n}}}{p^{n(2k-2-h)}}\right)(f)$
@@
Therefore the operator $\sum\limits_{\begin{subarray}{c}\alpha\geq 1\\ \gcd(\alpha,N)=1\end{subarray}}\frac{\mathbf{T}_{\alpha^{2}}}{\alpha^{2k-2-h}}$ is invertible, which implies
$\prod\limits_{p\mid N}\sum\limits_{n=1}^{\infty}\frac{\mathbf{T}_{p^{2n}}}{p^{n(2k-2-h)}}(f)=0.$
Let $n\coloneqq\prod\limits_{p\mid N}p$ be the radical of $N$. By [Ma24, Lem. 6.6], according to our assumption on $M$, there exists a sublattice $L\subseteq M$ such that $G_{L}\cong G_{M}\oplus G^{\prime}$, where $G^{\prime}=\left(\begin{array}[]{cc}0&\frac{1}{n}\\ \frac{1}{n}&0\end{array}\right)\bmod\mathbb{Z}$ under a basis $\{x_{1},x_{2}\}$ and $nx_{i}=0$. Then the cyclic subgroup $H=\langle x_{1}\rangle$ is an isotropic subgroup of $G_{L}$ and $H^{\perp}/H\cong G_{M}$.
@@
According to [Ma24, Lem. 6.3], also $\prod\limits_{p\mid N}\sum\limits_{n=1}^{\infty}\frac{\mathbf{T}_{p^{2n}}}{p^{n(2k-2-h)}}\uparrow_{H}(f)=0$. Let $v=v_{x_{1}+\gamma,x_{2}+\gamma,S}$ and then
@@
where we used Proposition 3.7 for the last equality. Since $\gamma$ is arbitrary, this implies that $f=0$. ∎
@@
There is a well-established criterion for determining whether $M\otimes\mathbb{Z}_{p}$ splits a hyperbolic plane, as mentioned in [Ni79]. For example, this occurs if ${\rm rank}_{p}(G_{M})<r-2$. In practical terms, this condition is not overly restrictive when $r$ is sufficiently large, as ${\rm rank}_{p}(G_{M})\leq r$ always holds.
Remark 4.7.
When $M\otimes\mathbb{Z}_{p}$ does not split a hyperbolic plane, ${\mathrm{Cusp}}_{k}(\rho_{M})$ may strictly contain ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M})$. Consider for example $M=L(N)$ where $L$ is unimodular of rank $m$ and $N$ is a square-free integer. For sufficiently large $m$ and $N$, there exist non-zero newforms $f$ of weight $m/2$ for $\Gamma_{0}(N)$.
The zero-component of any element in ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M})$ is an oldform. However, by [SV, Thm. 1.1], for any such newform $f$, the lift
$\sum_{g\in\Gamma_{0}(N)\backslash\mathrm{SL}_{2}(\mathbb{Z})}(f\cdot\mathfrak{e}_{0})|_{k,G_{M}}[g]$
has a non-zero newform as its zero-component, thus cannot lie in ${\mathrm{Cusp}}^{\theta}_{k}(\rho_{M})$.
5. Heegner divisors on orthogonal Shimura varieties
In this section, we review the general theory of Heegner divisors on orthogonal Shimura varieties and their connections with modular forms. We develop this correspondence to the Baily–Borel compactifications.
5.1. Orthogonal Shimura variety and its compactification